# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v1

@@ 4.1 Data Generation
We select the 36 anthropometric measurements for our models based on the selections of AnthroNet (Picetti et al., 2023) and an anthropometry study of the U.S. army (Gordon et al., 2014). They can be categorized into 23 lengths and 13 circumferences. Apart from the bone lenghts like arm length, thigh length, etc., this includes also detailed measurements like shoulder width, front torso height, lateral neck length, waist circumference, calf circumference, etc. A visualization and precise description of the selected measurements can be found in the supplementary.
Many existing datasets provide a wide range of different poses, but most incorporate the same humans. In order to learn a conversion model from anthropometric measurements to $\beta$ parameters, we need a lot of samples for different humans, no matter the pose. With given shape parameters, we can use the B2A function to compute the anthropometric measurements. Recall, B2A is a deterministic function measuring the anthropometric values from the meshes.
Because of the need for many different body shapes for the learning process, we use the AGORA (Patel et al., 2021) dataset. It consists of 1447 male and 1588 female subjects, which is comparably large. We are not able to use the larger dataset from AnthroNet (Picetti et al., 2023), since it uses its own mesh model and the authors did not publish their conversion to the SMPL-X model, which we want to use as it is most commonly used in research. Although comparably large, 1447/1588 subjects is still a very little amount of data to learn a model. Hence, we analyze the $\beta$ parameters in the AGORA dataset with the aim to randomly sample more data with realistic body shapes. Histograms (see Figure 2) of the occurring $\beta$ parameters show that their distribution roughly follows a normal distribution. Therefore, we train our models with randomly sampled data according to the distributions, either assuming a normal distribution fitted to the histograms or a uniform distribution with the same minimum and maximum values as in the data analysis. This means that we sample each $\beta$ parameter according to the selected distribution, create the mesh according to the sampled values and derive the anthropometric measurements with B2A. With this strategy, we can create a dataset with as many subjects as we need. We expect the analyzed data to not cover the full range of human body shapes as AGORA is a synthetic dataset. Therefore, we also train with extended distributions, meaning that we increase the standard deviation $\sigma$ to $\alpha_{n}\sigma$ in the case of a normal distribution or stretch the interval by a factor $\alpha_{u}$ in case of a uniform distribution.
Figure 2: Histograms and fitted normal distribution (orange) for the first two $\beta$ parameters for all male (left) and female (right) subjects of the AGORA (Patel et al., 2021) dataset.
@@ 4.2 Models
We use the same number of $\beta$ parameters for each gender as used in the AGORA dataset, meaning $11$ for male, $10$ for female, and $16$ for neutral subjects. With $36$ anthropometric measurements as input values and $10-16$ output values for our A2B models, the dimensionality of the data is low. Therefore, we experiment with Support Vector Regression (SVR) and with small neural networks (NN). We split the AGORA dataset in a 80% train, 15% test and 5% validation subset. For SVR, we additionally randomly sample $10,000$ subjects for training. We use a hyperparameter search based on the validation split to determine the optimal settings, which leads us to a radial basis function kernel, an error margin of $\epsilon=0.012$ and a regularization constant of $C=3791$. For the NNs, we randomly sample new data in each iteration. The hyperparameter search for the NNs results in a model with $4$ layers, $330$ neurons per layer, tanh as activation function, and Xavier Gloriot as initialization. We use mean squared error on the model output (the $\beta$ parameters) to train the models.
@@ 4.3 Results
Table 2: Results of our A2B models on the test split of the AGORA dataset. The first block ($\beta$) shows the error if we take the GT $\beta$ parameters, derive 36 anthropometric measurements (B2A), input them into the A2B models and evaluate the MSE of the predicted $\beta$ parameters in the scale $10^{-3}$. The second block ($A$) calculates B2A from the predicted $\beta$ parameters and evaluates the mean difference between the GT and predicted anthropometric measurements (all 36) in mm. Results are given for m(ale), f(emale), and n(eutral) models. A visualization is provided in the supplementary.
$\beta$ $A$
Model train data m f n m f n
@@
NN normal 2.62 4.34 18.0 0.356 0.392 1.711
NN normal ext. 1.87 3.69 14.8 0.248 0.285 1.384
NN uniform 5.08 1.25 18.3 0.243 0.268 2.381
NN uniform ext. 1.61 3.20 16.3 0.274 0.419 1.774
@@
SVR normal 4.08 17.8 59.0 2.975 4.303 14.63
SVR normal ext. 0.210 4.60 6.27 0.280 1.090 2.211
SVR uniform 0.0396 0.0350 0.162 0.124 0.284 0.214
SVR uniform ext. 0.0252 0.0193 0.306 0.082 0.136 0.164
We train each model (NN and SVR) for each gender and with different dataset variants: We train solely on the AGORA train split, as well as on uniformly and normally distributed randomly sampled data according to the data analysis, and we further extend the range of the data as described in Section 4.1 with $\alpha_{n}=\alpha_{u}=1.5$. The results are displayed in Table 2. We evaluate the performance of our models in two ways. At first, we calculate the error of the predicted and GT $\beta$ parameters. Second, we calculate the mean deviation of the anthropometric measurements of the meshes from the predicted and GT $\beta$ parameters ($A$). Therefore, this evaluation can further be seen as a kind of cycle consistency evaluation of A2B (our learned model) and B2A (the deterministic measuring function). We provide a visualization of the evaluation process in the supplementary. The anthropometric error is our main metric as these values reflect the desired body shape given as an input by the user and are further interpretable. The $\beta$ parameters are somehow arbitrary in their scale. For all genders and SVR, using an extended uniformly sampled dataset works best. For the NNs, a uniformly sampled dataset works best for male (m) and female (f) genders and an extended normally sampled dataset for the neutral (n) meshes. The results for the neutral model are worse in general, especially in the case of the NNs, which might be due to the fact that the neutral model needs to express a more diverse range of body shapes. Furthermore, the SVR achieves better results for all genders. Thus, we use these models now for all datasets, without any finetuning or adaptation to specific datasets.
@@ 5 Leveraging A2B Model Results for HME
Now that we have trained the A2B models, we can use them to generate precise body shape parameters upfront and reuse them for every evaluation of a specific person. In the next section, we describe how the A2B results can be used to improve existing HME models (see Section 5.1). Further, we introduce a new approach to HME (see Section 5.2). We leverage the good performance of a sequence-based 2D-to-3D uplifting HPE model and convert the 3D (stick-figure) poses to human meshes with the help of the A2B models. With this new approach, we achieve superior results compared to existing HME models.
@@
We evaluate all models on ASPset (Nibali et al., 2021), which is a 3D human pose dataset. It consists of various different sports motion clips performed by different subjects, recorded from three camera perspectives. We evaluate on the test set, which contains two subjects and 30 videos for each subject. In the test set, only one camera perspective is public, so we evaluate on this perspective. Evaluating SMPL-X meshes for ASPset is non-trivial. Regressing standard SMPL-X joints from SMPL-X meshes is built-in, but for all other keypoint definitions it is necessary to define a custom regressor. Since there is no regressor available for ASPset, we create a custom SMPL-X regressor for the keypoints head, head top, neck, l./r. hip, and torso. For left ankle and elbow, we mirror the corresponding right regressor, as it is not the exact mirrored version in the standard regressor. We use the code from (Russo, 2020) and update it to our needs.
@@
We further evaluate on fit3D (Fieraru et al., 2021), since this is the only sports dataset with public SMPL-X annotations. We evaluate on the SMPL-X joints here since they are available. We select a subset of 37 SMPL-X joints. Since our focus is mainly on the body and not on the hands and face, we remove a lot of these joints. A list of the selected joints can be found in the supplementary. For both datasets, we do not have access to the athletes to measure their anthropometric values. Therefore, we use the ground truth to obtain the anthropometric measurements that are needed as an input for the A2B models, details can be found in the supplementary. Since there is no GT available for the official test set evaluation on the evaluation server of fit3d, we split the official training dataset into a training, validation, and test set for our evaluations. We choose the last subject (s11) with all 188 videos as the test set.
@@
Sports datasets differ from most commonly used everyday activity datasets in the aspect that the poses are more diverse and the motions are faster, which makes sports datasets more difficult. In some cases, the poses are so difficult that some models do not detect a human at all. This makes a fair evaluation hard, since the standard MPJPE metric takes the mean of the joint position errors. Assuming a default pose for all frames where no person is detected would result in a very high error that shifts the mean enormously. Hence, we report the MPJPE only on the frames where persons are detected. Since mostly difficult frames are omitted, this will result in a slightly easier setting for methods that find less persons, but we include the number of missing frames in our results for comparison.
@@ 5.1 Improving HME Model Results
As already described, a major problem for HME based analyses is a varying body shape within a single video. Using $\beta$ parameters generated with A2B models solves this problem. The necessary 36 anthropometric measurements are either measured from the human directly or - as in our case - averaged from the provided GT (fit3D) or averaged from IK applied to the GT poses (ASPset). We call these measurements pseudo GT in the following. More details can be found in the supplementary. We combine existing HME models with the body shape estimated by our A2B models by replacing the estimated $\beta$ parameters with the ones predicted by the A2B models. We select three recent well performing models on the AGORA dataset (SMPLer-X (Cai et al., 2024), OSX (Lin et al., 2023), Multi-HMR (Baradel et al., 2024)), and the first HME method developed by the SMPL-X authors, SMPLify (Pavlakos et al., 2019). Since SMPLer-X is trained on the official training data of fit3D, an evaluation with this model is not meaningful and we omit it. Moreover, SMPLify-X is not SOTA any more and achieved the worst results for ASPset, therefore we omit it, too.
@@
The first evaluation contains the original result from the respective model, and evaluations where the pose from the model is kept, but the $\beta$ parameters are replaced with the A2B body shape parameters with pseudo GT input. In order to emphasize the variation of the body shape in the HME model results, we evaluate the standard deviation of the body height per person. Results are displayed in Table 3. We can see that for all models, replacing the estimated $\beta$ parameters by $\beta$ parameters from our A2B models with pseudo GT input leads to an improvement. For some models, the male meshes outperform the neutral ones and vice versa. The test sets of both datasets contain only male subjects, therefore we cannot evaluate female meshes here. Interestingly, the NN outperforms the SVR for all neutral experiments, although the SVRs achieved better results on the AGORA dataset evaluation. The reason could be that AGORA is a synthetic dataset and does not reflect reality. SMPLer-X achieves the best results for ASPset and Multi-HMR for fit3D, both with a significant margin. OSX performs worse on fit3D than on ASPset, but Multi-HMR performs better by a large margin and surpasses OSX. All methods benefit from our A2B $\beta$ parameters based on pseudo GT with improvements from 11 mm to 2 mm.
@@
Table 3: MPJPE results in mm for existing models on the test splits of ASPset (top) and fit3D (bottom). The second column (orig.) contains the original results, the other columns results with replaced $\beta$ parameters from our A2B models with pseudo GT anthropometric measurements as input and either male (m) or neutral (n) meshes, and the percentage of frames with no result (no r.). The $\sigma$ column displays the mean standard deviation of the body height per subject in cm for the original results. With A2B body shapes, $\sigma=0$.
Model orig. $\sigma$ NN m SVR m NN n SVR n no r. ($\downarrow$)
ASPset SMPLer-X (Cai et al., 2024) 86.02 1.8 78.88 78.50 78.38 78.47 0.11%
OSX (Lin et al., 2023) 92.34 0.2 89.58 89.28 89.44 89.56 0.10%
Multi-HMR (Baradel et al., 2024) 102.54 3.6 100.04 100.25 99.28 99.53 0.44%
SMPLify-X (Pavlakos et al., 2019) 138.18 13.0 127.73 127.36 126.78 126.89 0.02%
fit3D OSX (Lin et al., 2023) 99.20 4.8 93.22 92.58 91.72 92.21 3.45%
Multi-HMR (Baradel et al., 2024) 72.87 3.0 70.72 71.65 70.70 71.74 1.54%
@@
Table 4: MPJPE results in mm for existing models on the test split of the ASPset (top) and fit3D (bottom) datasets. The second column contains the original results, the other columns results with replaced $\beta$ parameters. Either the median $\beta$ parameters are used or the results from our A2B models with median anthropometric measurements from the respective model as input. For no r. see Table 3.
Model orig. median NN m SVR m NN n SVR n
ASPset SMPLer-X (Cai et al., 2024) 86.02 86.04 85.89 85.69 86.03 85.99
OSX (Lin et al., 2023) 92.34 92.37 92.44 92.17 92.33 92.38
Multi-HMR (Baradel et al., 2024) 102.54 102.05 102.59 102.96 102.07 102.16
SMPLify-X (Pavlakos et al., 2019) 138.18 133.59 133.81 133.46 133.56 133.48
fit3D OSX (Lin et al., 2023) 99.20 96.70 98.20 96.07 96.60 96.72
Multi-HMR (Baradel et al., 2024) 72.87 72.08 72.62 72.37 72.06 71.92
@@
We further evaluate the capabilities of a fixed body shape without available GT anthropometric measurements to ensure consistent body shapes also in the case that no measurements are available. Hence, we use the meshes originally created by the HME models, compute the anthropometric measurements of each mesh with B2A, calculate their median values and convert them to $\beta$ parameters via the A2B models. Results are displayed in Table 4. For SMPLer-X and OSX, using the median $\beta$ parameters lead to worse results on ASPset. Regarding ASPset, for all models except Multi-HMR, using our A2B models increases the performance slightly. Switching from the neutral output that these models all have to a gendered model works best in these cases, but the neutral A2B models also lead to a marginal improvement. Regarding fit3D, using the median $\beta$ parameters already enhances the results. Using $\beta$ parameters from an A2B model leads to further improvement, OSX achieves the best results with SVR and the male model, Multi-HMR with SVR and the neutral model.
@@
They can further be used to easily convert between neutral and gendered (male or female) models. In contrast, $\beta$ parameters are not transferable between models of different genders. Therefore, until now, the conversion could only be achieved by minimizing the vertex error between meshes of different genders in an iterative process. With our approach, we can now use the B2A fuction to obtain anthropometric measurements for a mesh of one gender and apply the A2B model of the other gender to these anthropometric measurements in order to get the corresponding $\beta$ parameters for this gender.
5.2 Human Mesh Estmation with Sequence Based 3D HPE and A2B Results

### v3

@@ 4.1 Data Generation
We select 36 anthropometric measurements for our models based on the selections of AnthroNet [28] and an anthropometry study of the U.S. army [14]. They can be categorized into 23 lengths and 13 circumferences. Apart from the bone lenghts like arm length, thigh length, etc., this includes also detailed measurements like shoulder width, front torso height, lateral neck length, waist circumference, calf circumference, etc. A visualization and precise description of all AMs can be found in the supplementary.
Many existing datasets provide a wide range of different poses, but most incorporate the same humans. For learning a conversion model from AMs to $\beta$ parameters, we need a lot of samples for different humans, no matter the pose. With given shape parameters, we can use the B2A function to compute the AMs. Recall, B2A is a deterministic function measuring the AMs from meshes in T-pose.
Because many different body shapes are required for the learning process, we use the AGORA [26] dataset. It consists of 1447 male and 1588 female subjects. We are not able to use the larger dataset from AnthroNet [28], since it uses its own mesh model and the authors did not publish their conversion to the SMPL-X model, which we want to use as it is most commonly used in research. Although comparably large, 1447/1588 subjects is still a little amount of data to learn a model. Hence, we analyze the $\beta$ parameters in the AGORA dataset with the aim to randomly sample more data with realistic body shapes. Histograms (see Figure 2) of the occurring $\beta$ parameters show that their distribution roughly follows a normal distribution. Therefore, we train our models with randomly sampled data according to these distributions, either assuming a normal distribution fitted to the histograms or a uniform distribution with the same minimum and maximum values as in the data analysis. This means that we sample each $\beta$ parameter according to the selected distribution, create the mesh according to the sampled values and derive the AMs with B2A. With this strategy, we can create a dataset with as many subjects as we need. As we do not expect the analyzed AGORA data to cover the full range of human body shapes, we also train with extended distributions, meaning that we increase the standard deviation $\sigma$ to $\alpha_{n}\sigma$ in the case of a normal distribution or stretch the interval by a factor $\alpha_{u}$ in case of a uniform distribution.
Figure 2: Histograms and fitted normal distribution (orange) for the first two $\beta$ parameters for all male (left) and female (right) subjects of the AGORA [26] dataset.
@@ 4.2 Models
We use the same number of $\beta$ parameters for each gender as used in the AGORA dataset, meaning $11$ for male, $10$ for female, and $16$ for neutral subjects. With $36$ AMs as input values and $10-16$ output values for our A2B models, the dimensionality of the data is low. Therefore, we experiment with Support Vector Regression (SVR) and with small neural networks (NN). We split the AGORA dataset in a 80% train, 15% test and 5% validation subset. For SVR, we additionally randomly sample $10,000$ subjects for training. We use a hyperparameter search based on the validation split to determine the optimal settings, which leads us to a radial basis function kernel, an error margin of $\epsilon=0.012$ and a regularization constant of $C=3791$. For the NNs, we randomly sample new data in each iteration. The hyperparameter search for the NNs results in a model with $4$ layers, $330$ neurons per layer, tanh as activation function, and Xavier Gloriot as initialization. We use mean squared error on the model output (the $\beta$ parameters) as training loss.
@@ 4.3 Results
[…]
train data m f n m f n
@@
NN norm. 2.62 4.34 18.0 0.356 0.392 1.711
NN norm. ext. 1.87 3.69 14.8 0.248 0.285 1.384
NN unif. 5.08 1.25 18.3 0.243 0.268 2.381
NN unif. ext. 1.61 3.20 16.3 0.274 0.419 1.774
@@
SVR norm. 4.08 17.8 59.0 2.975 4.303 14.63
SVR norm. ext. 0.210 4.60 6.27 0.280 1.090 2.211
SVR unif. 0.0396 0.0350 0.162 0.124 0.284 0.214
SVR unif. ext. 0.0252 0.0193 0.306 0.082 0.136 0.164
Table 2: Results of our A2B models on the test split of the AGORA dataset. The first block ($\beta$) shows the error if we take the GT $\beta$ parameters, derive 36 anthropometric measurements (B2A), input them into the A2B models and evaluate the MSE of the predicted $\beta$ parameters in the scale $10^{-3}$. The second block ($A$) calculates B2A from the predicted $\beta$ parameters and evaluates the mean difference between the GT and predicted AMs (all 36) in mm. Results are given for m(ale), f(emale), and n(eutral) models. Further visualizations are in the supplementary.
We train each model (NN and SVR) for each gender and with different dataset variants: We train solely on the AGORA train split, as well as on uniformly and normally distributed randomly sampled data according to the data analysis, and we further extend the range of the data as described in Section 4.1 with $\alpha_{n}=\alpha_{u}=1.5$. The results are displayed in Table 2. We evaluate the performance of our models in two ways. At first, we calculate the error of the predicted and GT $\beta$ parameters. Second, we calculate the mean deviation of the AMs of the meshes from the predicted and GT $\beta$ parameters ($A$). Therefore, this evaluation can further be seen as a kind of cycle consistency evaluation of A2B (our learned model) and B2A (the deterministic measuring function). We provide a visualization of the evaluation process in the supplementary. The anthropometric error is our main metric as these values reflect the desired body shape given as an input by the user and are further interpretable. The $\beta$ parameters are somehow arbitrary in their scale. For all genders and SVR, using an extended uniformly sampled dataset works best. For the NNs, a uniformly sampled dataset works best for male (m) and female (f) genders and an extended normally sampled dataset for the neutral (n) meshes. The results for the neutral model are worse in general, especially in the case of the NNs, which might be due to the fact that the neutral model needs to express a more diverse range of body shapes. Furthermore, the SVR achieves better results for all genders. Thus, we use these models for all datasets, without any finetuning or adaptation to specific datasets.
@@ 5 Leveraging A2B Model Results for HME
Now that we have trained the A2B models, we can use them to generate precise body shape parameters upfront and reuse them for every evaluation of a specific person. In the next section, we describe how the A2B results can be used to improve existing HME models (see Section 5.1). Further, we introduce a new approach to HME (see Section 5.2). We leverage the good performance of a sequence-based 2D-to-3D uplifting HPE model and convert the 3D stick-figure poses to human meshes with the help of our A2B models. With this approach, we achieve superior results compared to existing HME models. However, we want to emphasize that our approach is not exactly comparable to existing ones since it uses the additional information of anthropometric measurements. Since the performance of existing HME approaches is not good enough to be used for sports analyses, our main goal is to achieve the best possible performance with marginal additional information. As professional athletes are measured anyway, this results in actual no overhead in these usecases.
@@
We evaluate all models on the ASPset [25] 3D human pose dataset. It consists of various different sports motion clips performed by different subjects, recorded from three camera perspectives. We evaluate on the test set, which contains two subjects and 30 videos for each subject. In the test set, only one camera perspective is public, so we evaluate on this perspective. Evaluating SMPL-X meshes for ASPset is non-trivial. Regressing standard SMPL-X joints from SMPL-X meshes is built-in, but for all other keypoint definitions it is necessary to define a custom regressor. Since there is no regressor available for ASPset, we create a custom SMPL-X regressor [31].
@@
We further evaluate on fit3D [12], since this is the only sports dataset with public SMPL-X annotations. We evaluate the meshes and the SMPL-X joints since they are available. We select a subset of 37 SMPL-X joints. Since our focus is mainly on the body and not on the hands and face, we remove a lot of these joints and consider only the main body pose for MVE calculation. A list of the selected joints can be found in the supplementary. Hence, we achieve a fair comparison with this evaluation scheme. For both datasets, we do not have access to the athletes to measure them. Therefore, we simulate athlete measuring by measuring the GT meshes. Details can be found in the supplementary. Since there is no GT available for the official test set evaluation on the evaluation server of fit3d, we split the official training dataset into a training, validation, and test set for our evaluations. We perform a leave-one-out cross validation and average the results.
@@
Sports datasets differ from most commonly used everyday activity datasets in the aspect that the poses are more diverse and the motions are faster, which makes sports datasets more difficult. In some cases, the poses are so difficult that some models do not detect a human at all. This makes a fair evaluation hard, since the standard MPJPE metric takes the mean of the joint position errors. Assuming a default pose for all frames where no person is detected would result in a very high error that shifts the mean enormously. Hence, we report the MPJPE only on the frames where persons are detected. Since mostly difficult frames are omitted, this will result in a slightly easier setting for methods that find fewer persons, but we include the number of missing frames in our results for comparison.
@@ 5.1 Improving HME Model Results
A major problem for HME based analyses is a varying basic body shape within a single video. Existing HME models output different $\beta$ parameters for each frame. Exemplarily, we show the standard deviation of the body height of one subject in Table 3. Recall that these measurements and $\beta$ parameters are based on a T-pose mesh, hence varying poses have no influence on measuring and $\beta$ parameters. Using $\beta$ parameters generated with A2B models solves this problem. The necessary 36 measurements are either measured from the human directly, or averaged from the provided GT mesh (fit3D) or IK applied to the GT poses (ASPset). We call these measurements pseudo GT and include more details in the supplementary. We choose this process to simulate real measurements which exist for most professional athletes.We combine existing HME models with the body shape estimated by our A2B models by replacing the estimated $\beta$ parameters with the ones predicted by the A2B models. We select three recent well performing models on the AGORA dataset (SMPLer-X [3], OSX [19], Multi-HMR [1]), and the first HME method developed by the SMPL-X authors, SMPLify [27]. Since SMPLer-X is trained on the official training data of fit3D, an evaluation with this model is not meaningful and we omit it here. Moreover, SMPLify-X is not SOTA any more and achieved the worst results for ASPset. Therefore we omit it, too.
@@
The first evaluation contains the original result from the respective model, and evaluations where the pose from the model is kept, but the $\beta$ parameters are replaced with the A2B body shape parameters with pseudo GT input. Results are displayed in Table 3. The results for the MVE of the meshes for fit3D are included in Table 5. We can see that for all models, replacing the estimated $\beta$ parameters by $\beta$ parameters from our A2B models with pseudo GT input leads to an improvement. For one model, the gendered meshes outperform the neutral ones and for all other models, the neutral meshes perform best. We use the correct gender (male or female) of the subject in the gendered results. Interestingly, the NN outperforms the SVR for all neutral experiments, although the SVRs achieved better results on the AGORA dataset evaluation. The reason could be that AGORA is a synthetic dataset and does not reflect reality. SMPLer-X achieves the best results for ASPset and Multi-HMR for fit3D, both with a significant margin. OSX performs worse on fit3D than on ASPset, but Multi-HMR performs better by a large margin and surpasses OSX. All methods benefit from our A2B $\beta$ parameters based on pseudo GT with MPJPE improvements from 11 mm to 3 mm regarding both datasets and MVE improvements of approx. 8 mm regarding fit3D.
Model orig. $\sigma$ NN g SVR g NN n SVR n no r. $\downarrow$
ASPset SMPLer-X 86.0 2.9 78.9 78.5 78.4 78.5 0.11%
OSX 92.3 0.2 89.6 89.3 89.4 89.6 0.10%
Multi-HMR 102.5 3.6 100.0 100.3 99.3 99.5 0.44%
SMPLify-X 138.2 13.0 127.7 127.4 126.8 126.9 0.02%
fit3D OSX 94.2 3.9 88.9 88.6 87.1 87.2 3.45%
Multi-HMR 74.6 3.3 69.6 69.6 68.0 68.4 1.54%
@@
Table 3: MPJPE results in mm for existing models on the test splits of ASPset (top) and fit3D (bottom). The second column (orig.) contains the original results, the other columns results with replaced $\beta$ parameters from our A2B models with pseudo GT anthropometric measurements as input and either gendered (g) or neutral (n) meshes, and the percentage of frames with no result (no r.). The $\sigma$ column displays the mean standard deviation of the body height per subject in cm for the original results, while all A2B body shapes have $\sigma=0$.
Model orig. median NN g SVR g NN n SVR n
ASPset SMPLer-X 86.0 86.0 85.9 85.7 86.0 86.0
OSX 92.3 92.4 92.4 92.2 92.3 92.4
Multi-HMR 102.5 102.1 102.6 103.0 102.1 102.2
SMPLify-X 138.2 133.6 133.8 133.5 133.6 133.5
fit3D OSX 94.2 93.0 95.0 94.8 93.2 93.0
Multi-HMR 74.6 73.9 75.8 76.1 73.9 74.1
@@
Table 4: MPJPE results in mm for existing models on the test split of the ASPset (top) and fit3D (bottom) datasets. The second column contains the original results, the other columns results with replaced $\beta$ parameters. Either the median $\beta$ parameters are used or the results from our A2B models with median anthropometric measurements from the respective model as input.
@@
Although it is not our main goal, we further evaluate the capabilities of a fixed body shape without available GT measurements to ensure consistent body shapes in the case that no measurments are available. The simplest approach is to use the median of the $\beta$ parameters across all frames of the respective model. However, the $\beta$ parameters have no real meaning. Therefore, we compare this approach to taking the median of the anthropometric measurments of the generated meshes and then converting them to $\beta$ parameters via the A2B models. Results are displayed in Table 4. For SMPLer-X and OSX, using the median $\beta$ parameters lead to equal or even worse results on ASPset. Regarding ASPset, using our A2B models increases the performance of all models slightly. Switching from the neutral output that these models all have to a gendered model works best in most of these cases, but the neutral A2B models also lead to a marginal improvement. Regarding fit3D, using the median $\beta$ parameters already enhances the MPJPE and MVE results. Using $\beta$ parameters from an A2B model leads to the same improvement for both metrics, OSX achieves the best results with SVR and the neutral model, Multi-HMR with NN and the neutral model.
Figure 3: Overview of our inference pipeline. The pose and shape parameters are obtained either from IK applied to UU results (Sec. 5.2.2) or from an HME model (Sec. 5.1). In real applications, the anthropometric measurements will be taken directly from the humans. For our evaluations, we use the GT shape parameters and further experiment with the shape parameters of the respective model (IK or HME).
@@
Anthropometric measurements can further be used to easily convert between neutral and gendered (male or female) models. In contrast, $\beta$ parameters are not transferable between models of different genders. Therefore, until now, the conversion could only be achieved by minimizing the MVE between meshes of different genders in an iterative process. We can now use the B2A fuction to obtain measurements for a mesh of one gender and apply the A2B model of the other gender to these anthropometric measurements in order to get the corresponding $\beta$ parameters for this gender.
5.2 HME with Sequence Based 3D HPE and A2B